# Multi-Subarray FD-RIS Enhanced Multi-user Wireless Networks: With Joint Distance-Angle Beamforming

### v1

@@ IV-D Analysis on computing complexity and convergence of the proposed algorithm
1: Initialize ($\Big{(}\{\mathbf{w}^{(0)}_{k}\}_{k=1}^{K},\boldsymbol{\tau}^{(0)},\mathbf{f}^{(0)}\Big{)}$, tolerance accuracy $\breve{\epsilon}$;
@@
3:  Calculate the $W_{k}^{(q)}$ and $u_{k}^{(q)}$ with ($\Big{(}\{\mathbf{w}^{(q)}_{k}\}_{k=1}^{K},\boldsymbol{\tau}^{(q)},$
@@
6:  Update $\boldsymbol{\tau}^{(q+1)}$ based on the solution solved by Algo-
@@
7:  Update the $W_{k}^{(q)}$ and $u_{k}^{(q)}$ with the $\boldsymbol{\tau}^{(q+1)}$.
@@
10: Output: the optimal ($\Big{(}\{\mathbf{w}_{k}\}_{k=1}^{K},\boldsymbol{\tau},\mathbf{f}\Big{)}$.
@@ V Numerical Simulations
To validate the effectiveness of the proposed multi-subarray FD-RIS and the proposed iterative algorithm, we consider the following baseline schemes:
@@
    RIS-assisted scheme: In this scheme, the conventional RIS comprising an equivalent number of elements as the FD-RIS will be employed to support the wireless communication networks.
@@
    where $R_{k}$ is re-written as $R_{k}=\log_{2}\left(\frac{\operatorname{Tr}(\widehat{\mathbf{W}}\tilde{\mathbf{F}})+\sigma_{k}^{2}}{\operatorname{Tr}(\widehat{\mathbf{W}}\tilde{\mathbf{F}}_{-k})+\sigma_{k}^{2}}\right),$ with $\mathbf{F}=\mathbf{H}_{\mathrm{BR}}^{H}\boldsymbol{\Theta}_{k}^{H}\tilde{\boldsymbol{\Theta}}^{H}\mathbf{h}_{\mathrm{r}k}^{*}(\mathbf{H}_{\mathrm{BR}}^{H}\boldsymbol{\Theta}_{k}^{H}\tilde{\boldsymbol{\Theta}}^{H}\mathbf{h}_{\mathrm{r}k}^{*})^{H}$, $\tilde{\mathbf{F}}=\mathbf{E}_{K}\otimes\mathbf{F}$, $\tilde{\mathbf{F}}_{-k}=(\mathbf{E}_{-k}\mathbf{E}_{-k}^{T})\otimes\mathbf{F}$. Then, the first-order Taylor expansion is utilized to transform the non-convex objective function into convex. Similarly, we can also handle the time-delay subproblem.
@@
[…]
@@
To verify the distance-angle modulation capability of the multi-subarray FD-RIS, the 3D and 2D patterns of the users’ receiving energy distribution with the assistance of FD-RIS are presented in Fig. 4. Specifically, it is observed that the designed FD-RIS demonstrates the capability to manage the incoming signals in both distance and angle dimensions, which can precisely direct the signal energy towards the specific positions of users. Furthermore, to emphasize the considerable benefits of the FD-RIS in signal modulation across distance dimensions, we consider a scenario where two users are positioned at equal angles but varying distances, e.g., User 1 and User 3. The presented pattern results demonstrate that FD-RIS can accurately transmit the desired signal to the user’s location even under scenarios where the users at same direction with different distances. This capability, which is unattainable for conventional RIS in far-field communication scenarios, highlights the unique advantages of the FD-RIS.
Figure 5: Weighted sum rate versus the number of subarrays $L$ considering different number of antennas and power budget at the BS, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$.
Fig.5 gives the simulation results of the weighted sum rate versus the number of subarrays $L$ considering different number of antennas and power budget at the BS. In particular, we can find that the weighted sum rate shows a consistent gradual increase as the subarray size expands across all scenarios, but the increasing rates are decreasing. This is due to the fact that although the growing quantity of subarrays can enhance the level of control over incoming signals by providing additional degrees of freedom (DoFs), the system’s performance may still be restricted by other system configurations. Additionally, the SDR solution falls slightly behind the proposed approach in terms of performance gains obtained under the same scenario, yet it requires a significantly more computational resources to address optimization problems. We can also find that the performance improvement of the proposed scheme significantly exceeds that of the ZF scheme. Even when using fewer antennas ($N_{\mathrm{t}}=$ 5) to implement the proposed scheme, it still achieves higher performance level. These results demonstrate that the proposed scheme possesses a significant advantage in improving system performance. Furthermore, it is observed that the disparity in performance between the suggested method and the conventional RIS-assisted approach becomes more pronounced as the number of subarrays increases. This is primarily due to the enhanced DoFs of the FD-RIS, allowing for a more efficient concentration of signal energy towards the users. Because of this feature, when $L=$ 16, the FD-RIS can attain a performance comparable to that of the RIS-aided scheme with $L=$ 28, resulting in a reduction of almost 75% in elements. These underscore the potentials of the FD-RIS in augmenting communication performance.
Then, we investigate the varied trend of the weighted sum rate versus the number of antennas $N_{\mathrm{t}}$ considering different $L$ and $P_{\mathrm{tmax}}$, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$, as depicted in Fig. 6. Specifically, an increasing trend in the weighted sum rate is observed as the number of antennas increases in all scenarios, which is due to the fact that the increased $N_{\mathrm{t}}$ can provide more spatial DoFs to control different users’ signal and reduce the interference among users’ signals. Generally, under the same system configuration, the traditional RIS-assisted scheme exhibits the lowest performance, while the proposed scheme demonstrates the most significant performance improvement. Moreover, the performance gap between the two schemes becomes increasingly obvious as the number of antennas increases. This further indicates the superiority of the FD-RIS in enhancing communication systems’ performance. Additionally, while the disparity in performance between the SDR scheme and the suggested scheme diminishes as the quantity of antennas grows, the SDR solution requires a significant amount of computational resources and time.
Figure 6: Weighted sum rate versus the number of antennas $N_{\mathrm{t}}$ considering different number of subarrays and power budget at the BS, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$.
@@
Figure 7: Weighted sum rate versus the power budget at the BS $P_{\mathrm{tmax}}$ considering different number of antennas and subarrays, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$.
@@
We further explore the influence of the power budget at the BS on the weighted sum rate taking into account of different $L$ and $N_{\mathrm{t}}$ and weight factor with $\omega_{k}=$ 0.25, $k\in\mathcal{K}$, as shown in Fig. 7. In this simulation, two system configurations ($L=16,N_{\mathrm{t}}=10$ and $L=12,N_{\mathrm{t}}=7$) are employed to execute the proposed scheme, whereas the benchmark schemes are conducted utilizing the better system configuration available ($L=16,N_{\mathrm{t}}=10$). According to the presented simulation results, we can observe that the nearly linear growth of the weighted sum rate is observed with the expansion of the power budget $P_{\mathrm{tmax}}$ in the FD-RIS-aided schemes, while the increasing rate in the weighted sum rate gradually diminishes in the traditional RIS-supported scheme. This is due to the fact that the distance and angle modulation capability of the FD-RIS allows for the concentration of signal energy towards users, offering a substantial advantage in improving energy efficiency of the wireless networks. Similarly, the proposed scheme consistently achieves the highest performance gain for the communication system, and the performance gap between the proposed scheme and the RIS-assisted scheme becomes increasingly noticeable as $P_{\mathrm{tmax}}$ increases.
Figure 8: Achievable rate of each user under two sets of weight factor.
The weight factors, as previously stated in Section III, play a crucial role in managing users’ priority and ensuring fairness across users. To illustrate this feature, Fig. 8 investigates the influence of the weight factors on the achievable rate. Specifically, when each user possesses the equal weight factor, i.e., $\omega_{k}=0.25,k\in\mathcal{K}$, it is observed that the data rate attained by each user is inversely proportional to their distance from the FD-RIS. Consequently, User 1 enjoys the highest rate owing to its proximity, whereas User 2 experiences the lowest rate as a result of being the farthest away. To ensure the rate fairness, we appropriately increase the weight factor of users further away from the FD-RIS. In the case of unequal weight, each user’s weight factor is revised as $[\omega_{1},\omega_{2},\omega_{3},\omega_{4}]=[0.2,0.35,0.25,0.2]$. According to the presented simulation results under the scenario of unequal weights, we can achieve a more balanced data rate distribution, with User 2 receiving the highest priority in the communication system.
VI Conclusion
@@
Considering that the existing FD-RIS systems struggle with managing the undesired harmonic signals and exploiting the diversity of frequency offsets. To address these issues, this paper proposes a multi-subarray FD-RIS framework, in which the RIS is divided into multiple subarrays, each operating with a distinct time-modulation frequency to achieve frequency offset diversity. Additionally, to suppress the undesired harmonic interference, a novel time-modulation technique is introduced to periodically adjust the phase shifts of each element. According to the proposed multi-subarray FD-RIS framework, we first analytically derive its signal processing model. Then, the proposed multi-subarray FD-RIS is applied into the multi-user wireless communication systems to evaluate its potentials. Specifically, we establish a non-convex optimization problem to maximize the weighted sum rate of all users through designing the active beamforming, time delays and modulation frequencies. To address this problem, a novel iterative algorithm based on MMSE, bisection search method, Riemannian conjugate gradient algorithm and GCMMA method is proposed to solve it with low computing complexity. Extensive simulations are implemented and the obtained results indicate that the proposed multi-subarray FD-RIS can significantly enhance the energy efficiency of communication systems and thereby increase the system capacity when compared to the baseline schemes, including traditional RIS-assisted solutions.

### v2

@@ IV-D Analysis on computing complexity and convergence of the proposed algorithm
1: Initialize $\Big(\{\mathbf{w}^{(0)}_{k}\}_{k=1}^{K},\boldsymbol{\vartheta}^{(0)},\mathbf{f}^{(0)}\Big)$, tolerance accuracy $\breve{\epsilon}$;
@@
3:  Calculate the $W_{k}^{(q)}$ and $u_{k}^{(q)}$ with $\Big(\{\mathbf{w}^{(q)}_{k}\}_{k=1}^{K},\boldsymbol{\vartheta}^{(q)},$
@@
6:  Update $\boldsymbol{\vartheta}^{(q+1)}$ based on the solution solved by Algo-
@@
7:  Update the $W_{k}^{(q)}$ and $u_{k}^{(q)}$ with the $\boldsymbol{\vartheta}^{(q+1)}$.
@@
10: Output: the optimal $\Big(\{\mathbf{w}_{k}\}_{k=1}^{K},\boldsymbol{\tau},\mathbf{f}\Big)$.
@@
Figure 4: Three communication scenarios with various spatial correlations among users.
@@ V Numerical Simulations
In this section, the numerical simulations are implemented considering three scenarios featuring various spatial correlations among users to quantify the beamforming potential of the proposed multi-subarray FD-RIS and validate the effectiveness of the proposed iterative algorithm. Note that the number of users is set as $K=$ 4. The specific details of these scenarios are provided as follows:
Case 1: In this case, the spatial correlation among users is assumed to be negligible, which is a common assumption in conventional RIS-assisted communication systems. This setting corresponds to Case 1 illustrated in Fig. 4. In particular, all users’ positions are set as $(\theta_{\mathrm{r}1},\theta_{\mathrm{r}2},\theta_{\mathrm{r}3},\theta_{\mathrm{r}4})=(100^{\circ},110^{\circ},100^{\circ},120^{\circ})$, $(\phi_{\mathrm{r}1},\phi_{\mathrm{r}2},\phi_{\mathrm{r}3},\phi_{\mathrm{r}4})=(30^{\circ},70^{\circ},30^{\circ},150^{\circ})$, $(d_{\mathrm{r}1},d_{\mathrm{r}2},d_{\mathrm{r}3},d_{\mathrm{r}4})$
=(40m,75m,55m,40m)=(40m,75m,55m,40m).
Case 2: In this scenario, we assume that there is a strong spatial correlation between Bob 1 and Bob 3 users. This setting corresponds to Case 2 illustrated in Fig. 4. All users’ positions are set as $(\theta_{\mathrm{r}1},\theta_{\mathrm{r}2},\theta_{\mathrm{r}3},\theta_{\mathrm{r}4})=(100^{\circ},110^{\circ},100^{\circ},120^{\circ})$, $(\phi_{\mathrm{r}1},\phi_{\mathrm{r}2},\phi_{\mathrm{r}3},\phi_{\mathrm{r}4})=(20^{\circ},70^{\circ},30^{\circ},150^{\circ})$, $(d_{\mathrm{r}1},d_{\mathrm{r}2},d_{\mathrm{r}3},d_{\mathrm{r}4})$
=(40m,75m,55m,40m)=(40m,75m,55m,40m).
Case 3: It is assumed that Bob 1 and Bob 3 share the same spatial direction but are located at different distances, as illustrated in Case 3 of Fig. 4. It is worth noting that, in this extreme scenario, the beamforming capability of a conventional RIS fails to effectively manage the mutual interference between Bob 1 and Bob 3. All users’ positions are set as $(\theta_{\mathrm{r}1},\theta_{\mathrm{r}2},\theta_{\mathrm{r}3},\theta_{\mathrm{r}4})=(100^{\circ},110^{\circ},90^{\circ},120)$, $(\phi_{\mathrm{r}1},\phi_{\mathrm{r}2},\phi_{\mathrm{r}3},\phi_{\mathrm{r}4})=(120^{\circ},70^{\circ},30^{\circ},150^{\circ})$, $(d_{\mathrm{r}1},d_{\mathrm{r}2},d_{\mathrm{r}3},d_{\mathrm{r}4})$ =(40m,75m,55m,40m)=(40m,75m,55m,40m).
Additionally, we consider the following baseline schemes:
@@
    RIS-assisted scheme: In this scheme, the conventional RIS comprising an equivalent number of elements as the FD-RIS will be employed to support the wireless communication networks. It is worth noting that in this scheme, both the active beamforming at the BS and the reflection coefficients of the RIS need to be jointly designed. For the active beamforming variables, the same method presented in Section IV-A is adopted to efficiently solve the problem. As for the passive beamforming variables, the Riemannian conjugate gradient algorithm is employed to achieve effective optimization.
@@
     $~\text{ s.t. }\operatorname{Tr}(\widehat{\mathbf{W}})\leq P_{\mathrm{tmax}},$ (46a)
    $\qquad~\operatorname{rank}(\widehat{\mathbf{W}})=1,$ (46b)
    where $R_{k}$ is re-written as $R_{k}=\log_{2}\left(\frac{\operatorname{Tr}(\widehat{\mathbf{W}}\tilde{\mathbf{F}})+\sigma_{k}^{2}}{\operatorname{Tr}(\widehat{\mathbf{W}}\tilde{\mathbf{F}}_{-k})+\sigma_{k}^{2}}\right)$ with $\mathbf{F}=\mathbf{H}_{\mathrm{BR}}^{H}\boldsymbol{\Theta}_{k}^{H}\tilde{\boldsymbol{\Theta}}^{H}\mathbf{h}_{\mathrm{r}k}^{*}(\mathbf{H}_{\mathrm{BR}}^{H}\boldsymbol{\Theta}_{k}^{H}\tilde{\boldsymbol{\Theta}}^{H}\mathbf{h}_{\mathrm{r}k}^{*})^{H}$, $\tilde{\mathbf{F}}=\mathbf{E}_{K}\otimes\mathbf{F}$, $\tilde{\mathbf{F}}_{-k}=(\mathbf{E}_{-k}\mathbf{E}_{-k}^{T})\otimes\mathbf{F}$, $\mathbf{E}_{K}$ denotes a $K\times K$ identity matrix, $\mathbf{E}_{-k}$ is obtained by deleting the $k$-th column of $\mathbf{E}_{K}$. Then, the first-order Taylor expansion is utilized to transform the non-convex objective function into convex. The rank-one constraint is equivalently reformulated and incorporated into the objective function as a penalty term. By iteratively enforcing the penalty term towards zero, the rank-one will be met. The detailed process refer to [13]. Similarly, we can also handle the time-delay subproblem as
    $\max_{\mathbf{U}}~\sum_{k=1}^{K}\omega_{k}R_{k},$
     $~\text{ s.t. }\operatorname{diag}(\mathbf{U})=\mathbf{I}_{I\times 1},$ (47a)
    $\qquad~\operatorname{rank}(\mathbf{U})=1,$ (47b)
@@
    $\mathbf{H}_{\mathrm{r}k}^{H},$ $\tilde{\mathbf{G}}_{k}=\mathbf{H}_{\mathrm{r}k}\boldsymbol{\Theta}_{k}\mathbf{H}_{\mathrm{BR}}\sum_{j\neq k}^{K}\mathbf{w}_{j}\mathbf{w}_{j}^{H}\mathbf{H}_{\mathrm{BR}}^{H}\boldsymbol{\Theta}_{k}^{H}\mathbf{H}_{\mathrm{r}k}^{H}$, $\mathbf{H}_{\mathrm{r}k}=\operatorname{Diag}(\mathbf{h}_{\mathrm{r}k})$. The same approach is employed to address this optimization problem as in the active beamforming subproblem above.
@@
The system parameter values of executing the simulations are set as: the carrier frequency $f_{\mathrm{c}}=28$ GHz, the FD-RIS position $(d_{\mathrm{br}},\theta_{\mathrm{br}},\varphi_{\mathrm{br}})$= $(100m,30^{\circ},120^{\circ})$, g=g=1, the Rician factor $\beta_{1}=\beta_{2}=10$ dB, $\sigma_{k}^{2}=-110$ dBm, $k\in\mathcal{K}$, the number of elements in each subarray $M\times N=2\times 2$, $f_{\max}=20$ MHz, $f_{\min}=0.2$ MHz.
@@
To clearly demonstrate and verify the powerful distance–angle modulation capability of the multi-subarray FD-RIS, the third scenario is adopted as an illustrative example, and the corresponding 3D and 2D patterns of the users’ received energy distribution assisted by the FD-RIS are shown in Fig. 7. It is noted that, to facilitate pattern visualization, the elevation angles of all users are assumed identical and fixed at $90^{\circ}$. Specifically, it is observed that the designed FD-RIS demonstrates the capability to manage the incoming signals in both distance and angle dimensions, which can precisely direct the signal energy towards the specific positions of users. Furthermore, to emphasize the considerable benefits of the FD-RIS in signal modulation across distance dimensions, we consider a scenario where two users are positioned at equal angles but varying distances, e.g., User 1 and User 3. The presented pattern results demonstrate that FD-RIS can accurately transmit the desired signal to the user’s location even under scenarios where the users at same direction with different distances. This capability, which is unattainable for conventional RIS in far-field communication scenarios, highlights the unique advantages of the FD-RIS.
@@
Figure 8: Weighted sum rate versus the number of subarrays $L$ considering $N_{\mathrm{t}}=$ 10 and power budget $P_{\mathrm{tmax}}=$ 40 dBm at the BS, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$.
Fig.8 gives the simulation results of the weighted sum rate versus the number of subarrays $L$ with $N_{\mathrm{t}}=$ 10 and power budget $P_{\mathrm{tmax}}=$ 40 dBm at the BS, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$, under the different communication cases. In particular, we can find that the weighted sum rate shows a consistent gradual increase as the subarray size expands across all scenarios, but the increasing rates are decreasing. This is due to the fact that although the growing quantity of subarrays can enhance the level of control over incoming signals by providing additional degrees of freedom (DoFs), the system’s performance may still be restricted by other system configurations. Additionally, the SDR solution exhibits slightly inferior performance compared to the proposed approach when $L$ ranges from 4 to less than 24. However, as $L$ exceeds 24, the SDR scheme attains superior performance gains, albeit at the expense of substantially increased computational complexity required to solve the associated optimization problems. We can also find that the performance improvement of the proposed scheme significantly exceeds that of the ZF scheme. These results demonstrate that the proposed algorithm possesses a significant advantage in optimization ability. Furthermore, it is observed that the disparity in performance between the FD-RIS-assisted scheme and the conventional RIS-assisted scheme becomes more pronounced as the number of subarrays increases. This is mainly because the FD-RIS provides higher DoFs, allowing signal energy to be more effectively focused on the users via distance–angle beamforming. According to the results presented in Case 1–Case 3, the performance gain provided by the RIS gradually decreases as the spatial correlation among users increases, especially when some users have the same spatial direction. This is because, as the spatial correlation among users gradually increases, the angle-only beamforming capability of conventional RIS is significantly weakened. In particular, when multiple users are located in the same or similar directions, the RIS can hardly distinguish signals among users through beam control in a single angular dimension, resulting in degraded beamforming performance. In such cases, the RIS can typically guarantee the QoS for only one user, while the QoS of other users becomes severely limited. In contrast, the FD-RIS-aided systems (including the FD-RIS, SDR, and ZF schemes) maintain nearly the same level of performance gain with the distance-angle beamforming, showcasing the high adaptability and robustness for transmission environment of FD-RIS.
Then, we investigate the varied trend of the weighted sum rate versus the number of antennas $N_{\mathrm{t}}$ considering different $L$ and $P_{\mathrm{tmax}}$, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$, under three different communication scenarios, as depicted in Fig. 9. Specifically, an increasing trend in the weighted sum rate is observed as the number of antennas increases in all scenarios, which is due to the fact that the increased $N_{\mathrm{t}}$ can provide more spatial DoFs to transmit different users’ desired signal and effectively manage the interference among users’ signals. Next, a more detailed analysis of Fig. 9 is presented. Specifically, according to the results shown in Case 1, it can be observed that in scenarios with low spatial correlation among users, when $L$ is small, both the FD-RIS and the conventional RIS bring almost identical performance gains to the system. In this case, since the channels of different users are nearly independent, the beamforming capability of the conventional RIS is already sufficient to effectively enhance the desired signals and manage inter-user interference, meaning the additional dimensional advantage of the FD-RIS has not yet been fully utilized. However, as $L$ (for example, $L=$ 25), the FD-RIS-assisted system exhibits only a slight advantage over the conventional RIS in terms of weighted sum rate. This further confirms that under low spatial correlation, the system performance improvement of RIS is mainly constrained by the number of physical resources rather than the dimensionality of beamforming. When the spatial correlation increases, the performance advantage of FD-RIS becomes increasingly significant. Specifically, as observed in Case 2 and Case 3, the FD-RIS maintains almost the same performance gain as in Case 1. In contrast, the conventional RIS, which only supports one-dimensional angular beamforming, fails to flexibly control signal energy and manage interference. As a result, its overall system performance deteriorates rapidly with increasing spatial correlation. This further validates the superior capability of FD-RIS’s two-dimension beamforming.
Figure 9: Weighted sum rate versus the number of antennas, $N_{\mathrm{t}}$, considering different number of subarrays and power budget at the BS, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$.
Figure 10: Weighted sum rate versus the power budget at the BS, $P_{\mathrm{tmax}}$, considering different number of antennas and subarrays, and $\omega_{k}=$ 0.25, $k\in\mathcal{K}$.
@@
To quantify the energy efficiency of FD-RIS and RIS, we further explore the influence of the power budget at the BS on the weighted sum rate taking into account of weight factor with $\omega_{k}=$ 0.25, $k\in\mathcal{K}$, and different $L$ and $N_{\mathrm{t}}$, as shown in Fig. 10. In this simulation, two system configurations ($L=16,N_{\mathrm{t}}=10$ and $L=12,N_{\mathrm{t}}=7$) are employed to execute the proposed and baseline schemes. According to the presented simulation results, we can observe that the nearly linear growth of the weighted sum rate is observed with the expansion of the power budget $P_{\mathrm{tmax}}$ in the FD-RIS-aided schemes, while the increasing rate in the weighted sum rate gradually diminishes in the traditional RIS-supported scheme, particularly in scenarios with higher spatial correlation (i.e., Case 2 and Case 3). This is because the distance-angle beamforming capability of the FD-RIS enables it to effectively concentrate signal energy toward the intended users while mitigating mutual interference among them, thereby providing a significant advantage in enhancing the energy efficiency of wireless networks. Similarly, the FD-RIS maintains almost the same performance gain under different levels of spatial correlation, presenting remarkable environmental adaptability and robustness.
VI Conclusion and Prospect
@@
Considering that the existing FD-RIS systems struggle with managing the undesired harmonic signals and exploiting the diversity of frequency offsets. To address these issues, this paper proposes a multi-subarray FD-RIS framework, in which the RIS is divided into multiple subarrays, each operating with a distinct time-modulation frequency to achieve frequency offset diversity. Additionally, to suppress the undesired harmonic interference, a novel time-modulation technique is introduced to periodically adjust the phase shifts of each element. According to the proposed multi-subarray FD-RIS framework, we first analytically derive its signal processing model. Then, the proposed multi-subarray FD-RIS is applied into the multi-user wireless communication systems to evaluate its potentials. Specifically, we establish a non-convex optimization problem to maximize the weighted sum rate of all users through designing the active beamforming, time delays and modulation frequencies. To address this problem, the MMSE method is first leveraged to transformed the objective function and then a novel iterative algorithm based on Lagrange multiplier method, bisection search method, Riemannian conjugate gradient algorithm and GCMMA method is proposed to solve these three subproblem with low computing complexity. To evaluate the performance of the proposed scheme, three communication scenarios with different levels of spatial correlation among users are considered. Extensive simulation results indicate that the proposed multi-subarray FD-RIS scheme, supported by the distance–angle beamforming mechanism, can stably maintain overall capacity under various scenarios, demonstrating excellent robustness and adaptability. Moreover, compared with conventional RIS, the FD-RIS significantly enhances system performance through its powerful beamforming capability, with the advantage being particularly pronounced when users share the same spatial angle.
This paper provides an initial contribution to the multi-subarray FD-RIS. For future works, several pertinent potential use cases can be recognized: (i) ISAC: Leveraging distance–angle beamforming of FD-RIS, the ISAC system can simultaneously obtain distance and angle information of targets based on the echo signals. (ii) Secure communication: The single-dimensional control of traditional RIS may lead to “security blind zone” in communication systems. Specifically, when an eavesdropper and a legitimate user locate at the same spatial direction, the system struggles to distinguish between them, making it difficult to effectively suppress eavesdropping attempts. In contrast, FD-RIS can exploit differences in the distance domain to adjust the signal propagation, thereby overcoming this challenge.